OpenStudy (anonymous):

How do you fine a gerneral form equation for (1, -3) and (5,-3)

hero (hero):

The general form for a linear equation is `Ax + By = C` But it would be easier if you began with the slope-intercept form: `y = mx + b`

hero (hero):

You can put it back in the general form later.

OpenStudy (anonymous):

how do you find b? because to find m is y2-y1/x2-x1

hero (hero):

Well, if you find m, you can find b because you can use one of the points and m in the equation y = mx + b

hero (hero):

If you substitute values for x,y, and m, the only thing that would be left to find is b.

hero (hero):

You can use basic algebra to solve for b

OpenStudy (anonymous):

ok so its -3 -3 / 5-1 which would be 0/4 soooo what is it?

hero (hero):

You're telling me you don't know what 0/4 equals?

OpenStudy (anonymous):

undefined! lol sorry so thats what i write?

hero (hero):

Only if zero is in the denominator would it be undefined.

OpenStudy (anonymous):

dang. ok so once i fined b how do i put it in gerneral form?

hero (hero):

Show me what your equation is so far.

OpenStudy (anonymous):

y=0/4x-12 right?

hero (hero):

Where did the -12 come from?

OpenStudy (anonymous):

i plugged in (5,-3) into y=mx+b so -3 = (0/4)(5)+b and solved for b